# Sharp error bounds for edge-element discretisations of the high-frequency Maxwell equations

### v2

@@ 7.3 $S=\psi(\eta^{-1}{\cal P})$ increases regularity
We now apply Lemma 7.15 with $u=(\eta^{-1}{\cal P})^{m}\psi(\eta^{-1}{\cal P})\Pi_{1}v$ for arbitrary $v\in{\cal V}$. The proof that this $u\in{\cal H}$ is very similar to the proof of Lemma 7.16, using Lemma 7.10 – the key points are that (i) any compactly supported function of $\eta^{-1}{\cal P}$ is bounded $\Pi_{1}{\cal V}\to\Pi_{1}{\cal V}$, and (ii) the ${\cal H}$ norm essentially just adds another power of $\eta^{-1}{\cal P}$ by the Gårding inequality (7.7)).
@@ 7.5 Quasi-optimality of $\Pi^{\#}_{h}$
$\big{(}1-C_{1}\gamma_{\rm dv}(P)\big{)}\left\|u-u_{h}\right\|_{{\cal H}}\leq C_{2}\Big{(}\left\|(I-\Pi_{h})u\right\|_{{\cal H}}+\left\|S\Pi_{1}(u-u_{h})\right\|_{{\cal V}}\Big{)}\quad\text{ for all }v\in{\cal H}.$
@@
The arguments at the start of Lemma 6.3 show that it is sufficient to prove the bound
@@
(note that (8.1) is identical to (6.4) apart from the $S$ multiplying $\Pi_{1}(u-u_{h})$ on the right-hand side).
By coercivity of $P^{\#}=P+\Pi^{*}_{1}S^{2}\Pi_{1}$ on $\Pi_{1}{\cal H}$ (7.14),
@@
$\qquad\leq C_{3}\bigg{(}1-C_{2}\left\|(I-\Pi_{h})(P^{\#})^{-1}\Pi_{1}^{*}\eta S\Pi_{1}\right\|_{{\cal V}\to{\cal H}}\big{\|}(I-\Pi^{\#}_{h})(P^{*})^{-1}\Pi_{1}^{*}\eta S\Pi_{1}\big{\|}_{{\cal V}\to{\cal H}}$
@@
By Lemma 7.18 the projection $\Pi^{\#}_{h}$ is well-defined and satisfies (7.32) if $\gamma_{\rm dv}(P^{*})$ is sufficiently small. Therefore, the instances of $(I-\Pi^{\#}_{h})$ in Lemma 8.3 can be replaced (up to constants) by $(I-\Pi_{h})$.
@@ 9 Recap of the regularity result of Weber [68]
The following result is [68, Theorem 2.2], where we observe that this result – originally proved for real-valued coefficients – immediately generalises to complex-valued coefficients. Recall the definitions of the piecewise spaces $H^{j}_{\rm pw}(\Omega)$ (2.13) and the associated norm $\|\cdot\|_{H^{j}_{{\rm pw},k}(\Omega)}$ (2.14).
@@
Then there exists $C>0$ such that, for all $10\leq\ell\leq\kappa-$ 1, if either $u\times n=0$ or $(\zeta u)\cdot n=0$ on $\partial\Omega$ then
@@
The result for $\zeta$ real-valued and symmetric positive-definite and with norms not weighted with $k$ is [68, Theorem 2.2]. Repeating the proof but now weighting each derivative by $k^{-1}$ gives the bound (9.1).
We now outline why the result holds for complex-valued $\zeta$ with $\operatorname{Re}\zeta\geq c>0$. The proof of [68, Theorem 2.2] begins by localising and mapping the boundary to a half-plane (using [68, Lemma 3.1]) – this is unaffected by the change in assumptions on $\zeta$. The parts of the proof that depend on $\zeta$ then involve
@@
    the decomposition of an arbitrary $L^{2}$ vector field $F$ into $F_{1}+F_{2}$, where $F_{1}=\nabla f$ and $\mathop{{\rm div}}(\zeta F_{2})=0$, and either $f\in H^{1}_{0}(\Omega)$ and $\zeta F_{2}\in H(\mathop{{\rm div}};\Omega)$ or $f\in H^{1}(\Omega)$ and $\zeta F_{2}\in H(\mathop{{\rm div}};\Omega)$ with $(\zeta F_{2})\cdot n=0$ on $\partial\Omega$ [68, Lemmas 3.4 and 3.5].
The arguments in Point 1 go through verbatim (noting that $\zeta$ is still invertible). The results in Point 2 are quoted from [67, Lemmas 3.8 and 3.9], where they are proved using projections in the $L^{2}(\Omega)$ inner product weighted with $\zeta$. When $\zeta$ is complex valued, the results in Point 2 can be proved via the following. For the first result (when $f\in H^{1}_{0}(\Omega)$), given $F$, let $f\in H^{1}_{0}(\Omega)$ be the solution of the variational problem
@@
When $\operatorname{Re}\zeta\geq c>0$, the solution of (9.3) (a Laplace Neumann problem) is unique up to constants, so that $F_{2}:=F-\nabla f$ is uniquely defined. Now (9.3) for $w\in H^{1}_{0}(\Omega)$ is the statement that $\mathop{{\rm div}}(\zeta F_{2})=0$ and (9.3) for $w\in H^{1}(\Omega)$ implies that $n\cdot(\zeta F)=0$ by, e.g., [56, Equation 3.33]. ∎
@@ 10.1 Curved tetrahedral mesh
We consider a partition of $\Omega$ into a conforming mesh ${\cal T}_{h}$ of (curved) tetrahedral elements $K$ as in, e.g., [52, Assumption 3.1]. For $K\in{\cal T}_{h}$ we denote by $\mathscr{F}_{K}:\widehat{K}\to K$ the mapping between the reference tetrahedra $\widehat{K}$ and the element $K$. We further assume that the mesh ${\cal T}_{h}$ is conforming with the partition $\{\Omega_{i}\}_{i=1}^{n}$ of $\Omega$ from Assumption 1.2. This means that for each $K\in{\cal T}_{h}$, there is a unique $i\in\{1,\ldots,n\}$ such that $K\subset\overline{\Omega_{i}}$.
@@ 10.2 Nédélec finite element space
$N_{p}(\widehat{K})=P_{p-1}(\widehat{K})+x\times P_{p-1}(\widehat{K}),$
where $P_{s}(\widehat{K})$ consists of functions such that each component is a polynomial of degree $\leq s$ defined over $\widehat{K}$. (Note that in [23, Chapter 15] the lowest-order elements correspond to $p=0$, whereas here they correspond to $p=1$.) The associated approximation space is obtained by mapping the Nédélec polynomial space to the mesh cells through a Piola mapping (see (B.2) below), leading to
@@
for $11\leq|\alpha|\leq p+$ 1, where $h_{K}$ is the diameter of $K$.
Note that the bound (10.1) with $|\alpha|=1$ corresponds to the mesh elements $K\in\mathcal{T}_{h}$ being shape regular (as in, e.g., [52, Equation 3.3]).
@@ 10.3 High-order interpolation
[…]
@@
$\|v-\mathcal{J}_{h}v\|_{L^{2}(\Omega)}\leq C\bigg{(}\frac{h}{L}\bigg{)}^{\ell}\sum_{i=1}^{n}\sum_{j=0}^{\ell}L^{j}\Big{(}|u|_{H^{j}(\Omega_{i})}+(kh_{K})|k^{-1}{\rm curl}\,u|_{H^{j}(\Omega_{i})}\Big{)}$
@@
$\|k^{-1}{\rm curl}\,(v-\mathcal{J}_{h}v)\|_{L^{2}(\Omega)}\leq C\bigg{(}\frac{h}{L}\bigg{)}^{\ell}\sum_{i=1}^{n}\sum_{j=1}^{\ell}L^{j}|{\rm curl}\,u|_{H^{j}(\Omega_{i})}.$
@@
Points 1, 2, and 3 are proved in §11.1, §11.2, and §11.3, respectively. Regarding Point 3: we show in Lemma 11.5 below that $\max\{\gamma_{\rm dv}(P),\gamma_{\rm dv}(P^{*})\}\leq Ckh(1+kh)$, i.e., for the first inequality in (2.24) to hold, $kh$ must be sufficiently small. This condition is indeed weaker when $kL\gg 1$ than the condition “$(kh)^{2p}C_{\rm sol}$ is sufficiently small” arising from the second inequality in (2.24), since $C_{\rm sol}\geq CkL$ (as recalled in Remark 1.6).
@@
Since $H_{0}({\rm curl}\,;\Omega)\supset C^{\infty}_{0}(\Omega)$ (by, e.g., [56, Equation 3.42]) and $C^{\infty}_{0}(\Omega)$ is dense in $L^{2}(\Omega)$, ${\cal H}=H_{0}({\rm curl}\,;\Omega)$ is dense in ${\cal V}=L^{2}(\Omega)$.
@@
With ${\cal H}=H_{0}({\rm curl}\,,\Omega)$, by, e.g., [56, Theorem 3.31],
@@
Let $w:=k^{-1}\zeta_{1}{\rm curl}\,v$ so that $k^{-1}{\rm curl}\,w=f$. Observe that $\mathop{{\rm div}}(\zeta_{1}^{-1}w)=0$ and $\mathop{{\rm div}}f=0$. With ${\rm div}_{T}$ the surface divergence on $\partial\Omega$, by, e.g., [56, Equation 3.52], on $\partial\Omega$,
@@
To prove that Part (ii) of Assumption 2.3 holds, we seek to apply Lemma 11.2 with $v=\Pi_{1}u$, $\zeta_{1}=\mu^{-1},$ and $\zeta_{2}=\mathcal{E}=\epsilon$. Since $C^{\infty}(\overline{D})$ is dense in $H^{s}(D)$ for all $s\in\mathbb{R}$ and $D\subset\mathbb{R}^{3}$ open (see, e.g., [49, Page 77]), $L^{2}(\Omega)$ is dense in $(H^{j}_{\rm pw}(\Omega))^{*}$ for $j\geq 1$, so the assumption that ${\cal V}$ is dense in $({\cal Z}^{j})^{*}=(Z^{j})^{*}$ for $j\geq 1$ is satisfied. The bound (2.10) then follows from (11.3) if $\Pi_{1}u$ is in $H_{0}({\rm curl}\,,\Omega)$ and satisfies $\mathop{{\rm div}}(\epsilon\Pi_{1}u)=0$. Recall from (2.8) that $\Pi_{0}:{\cal H}\to{\cal H}$, and thus also $\Pi_{1}:{\cal H}\to{\cal H}$. Thus, $u\in H_{0}({\rm curl}\,,\Omega)$ implies that $\Pi_{1}u\in H_{0}({\rm curl}\,,\Omega)$. For the zero-divergence condition, observe that, since gradients are always inside the kernel of ${\rm curl}\,$,
@@
Thus, $\mathop{{\rm div}}\Pi_{1}^{*}v=0$ for all $v\in L^{2}(\Omega)$ (by, e.g., [56, Equation 3.33]). By (3.2), $\mathop{{\rm div}}(\epsilon\Pi_{1}u)=-\mathop{{\rm div}}(\mathcal{E}\Pi_{1}u)=-\mathop{{\rm div}}(\Pi_{1}^{*}\mathcal{E}\Pi_{1}u)$ so that $\mathop{{\rm div}}(\epsilon\Pi_{1}u)=0$ as required.
@@
By, e.g., [56, Theorem 3.41], the kernel of the curl operator in $H_{0}({\rm curl}\,,\Omega)$ equals $(\nabla H^{1}_{0}(\Omega))\oplus K_{N}(\Omega)$, where
@@
(the normal cohomology space); note that the dimension of $K_{N}(\Omega)$ equals the number of connected components of $\partial\Omega$ minus one; see, e.g., [56, Theorem 3.42]. Let $\Pi^{{\cal V}}_{\nabla H^{1}_{0}(\Omega)}$ and $\Pi^{{\cal V}}_{K_{N}(\Omega)}$ be the ${\cal V}$-orthogonal projections onto $\nabla H^{1}_{0}(\Omega)$ and $K_{N}(\Omega)$, respectively, so that
@@
$(\nabla\phi,\nabla v)_{L^{2}(\Omega)}=(f,\nabla v)_{L^{2}(\Omega)}\quad\text{ for all }v\in H^{1}_{0}(\Omega).$ (11.12)
@@
We now apply Theorem 9.1 with $u=\nabla\phi$, $\zeta=I$, and $\kappa=m$ (note that this is allowed because $\Omega$ is $C^{m+1}$ with respect to the partition). Since $\phi\in H^{1}_{0}(\Omega)$ implies that $\nabla\phi\in H_{0}({\rm curl}\,;\Omega)$ (see, e.g., [56, Equation 3.60/§B.3]), the boundary condition on $u$ in Theorem 9.1 is satisfied and, for $\ell=0,\ldots,m-1$,
@@
$\leq C\Big{(}\left\|\nabla\phi\right\|_{L^{2}(\Omega)}+\left\|\nabla w\right\|_{H^{j-1}_{{\rm pw},k}(\Omega)}\Big{)}$
$\leq C\Big{(}\left\|\nabla\phi\right\|_{L^{2}(\Omega)}+\left\|\nabla w\right\|_{Z^{j}_{k}}\Big{)}=C\Big{(}\left\|\nabla\phi\right\|_{{\cal V}}+\left\|\nabla w\right\|_{{\cal Z}^{j}}\Big{)};$

### v3

@@ 7.3 $S=\psi(\eta^{-1}{\cal P})$ increases regularity
[…]
@@ 7.5 Quasi-optimality of $\Pi^{\#}_{h}$
$\big(1-C_{1}\gamma_{\rm dv}(P)\big)\left\|u-u_{h}\right\|_{{\cal H}}\leq C_{2}\Big(\left\|(I-\Pi_{h})u\right\|_{{\cal H}}+\left\|S\Pi_{1}(u-u_{h})\right\|_{{\cal V}}\Big)\quad\text{ for all }v\in{\cal H}.$ (8.1)
@@
We first argue exactly as at the start of Lemma 6.3. By the triangle inequality, (2.8), and (5.1),
i.e.,
$\big(1-C\gamma_{\rm dv}(P)\big)\left\|u-u_{h}\right\|_{{\cal H}}\leq C\left\|(I-\Pi_{h})u\right\|_{\cal H}+\left\|\Pi_{1}(u-u_{h})\right\|_{\cal H}.$ (8.2)
We now claim that it is sufficient to prove the bound
@@
(note that (8.3) is identical to (6.4) apart from the $S$ multiplying $\Pi_{1}(u-u_{h})$ on the right-hand side). Indeed, inputting (8.3) into (8.2) and using again (5.1), we find (8.1).
Now, by coercivity of $P^{\#}=P+\Pi^{*}_{1}S^{2}\Pi_{1}$ on $\Pi_{1}{\cal H}$ (7.14),
@@
$\qquad\leq C_{3}\bigg(1-C_{2}\big\|(I-\Pi_{h})(P^{\#})^{-1}\Pi_{1}^{*}\eta S\Pi_{1}\big\|_{{\cal V}\to{\cal H}}\big\|(I-\Pi^{\#}_{h})(P^{*})^{-1}\Pi_{1}^{*}\eta S\Pi_{1}\big\|_{{\cal V}\to{\cal H}}$
@@
[…]
@@ 9 Recap of the regularity result of Weber [69]
The following result is [69, Theorem 2.2], where we observe that this result – originally proved for real-valued coefficients – immediately generalises to complex-valued coefficients. Recall the definitions of the piecewise spaces $H^{j}_{\rm pw}(\Omega)$ (2.13) and the associated norm $\|\cdot\|_{H^{j}_{{\rm pw},k}(\Omega)}$ (2.14).
@@
Then there exists $C>0$ such that, for all $0\leq\ell\leq\kappa-1$, if either $u\times n=0$ or $(\zeta u)\cdot n=0$ on $\partial\Omega$ then
@@
The result for $\zeta$ real-valued and symmetric positive-definite and with norms not weighted with $k$ is [69, Theorem 2.2]. Repeating the proof but now weighting each derivative by $k^{-1}$ gives the bound (9.1).
We now outline why the result holds for complex-valued $\zeta$ with $\operatorname{Re}\zeta\geq c>0$. The proof of [69, Theorem 2.2] begins by localising and mapping the boundary to a half-plane (using [69, Lemma 3.1]) – this is unaffected by the change in assumptions on $\zeta$. The parts of the proof that depend on $\zeta$ then involve
@@
    the decomposition of an arbitrary $L^{2}$ vector field $F$ into $F_{1}+F_{2}$, where $F_{1}=\nabla f$ and $\mathop{{\rm div}}(\zeta F_{2})=0$, and either $f\in H^{1}_{0}(\Omega)$ and $\zeta F_{2}\in H(\mathop{{\rm div}};\Omega)$ or $f\in H^{1}(\Omega)$ and $\zeta F_{2}\in H(\mathop{{\rm div}};\Omega)$ with $(\zeta F_{2})\cdot n=0$ on $\partial\Omega$ [69, Lemmas 3.4 and 3.5].
The arguments in Point 1 go through verbatim (noting that $\zeta$ is still invertible). The results in Point 2 are quoted from [68, Lemmas 3.8 and 3.9], where they are proved using projections in the $L^{2}(\Omega)$ inner product weighted with $\zeta$. When $\zeta$ is complex valued, the results in Point 2 can be proved via the following. For the first result (when $f\in H^{1}_{0}(\Omega)$), given $F$, let $f\in H^{1}_{0}(\Omega)$ be the solution of the variational problem
@@
When $\operatorname{Re}\zeta\geq c>0$, the solution of (9.3) (a Laplace Neumann problem) is unique up to constants, so that $F_{2}:=F-\nabla f$ is uniquely defined. Now (9.3) for $w\in H^{1}_{0}(\Omega)$ is the statement that $\mathop{{\rm div}}(\zeta F_{2})=0$ and (9.3) for $w\in H^{1}(\Omega)$ implies that $n\cdot(\zeta F)=0$ by, e.g., [57, Equation 3.33]. ∎
@@ 10.1 Curved tetrahedral mesh
We consider a partition of $\Omega$ into a conforming mesh ${\cal T}_{h}$ of (curved) tetrahedral elements $K$ as in, e.g., [53, Assumption 3.1]. For $K\in{\cal T}_{h}$ we denote by $\mathscr{F}_{K}:\widehat{K}\to K$ the mapping between the reference tetrahedron $\widehat{K}$ and the element $K$. We further assume that the mesh ${\cal T}_{h}$ is conforming with the partition $\{\Omega_{i}\}_{i=1}^{n}$ of $\Omega$ from Assumption 1.2. This means that for each $K\in{\cal T}_{h}$, there is a unique $i\in\{1,\ldots,n\}$ such that $K\subset\overline{\Omega_{i}}$.
@@ 10.2 Nédélec finite element space
[…]
@@
for $1\leq|\alpha|\leq p+1$, where $h_{K}$ is the diameter of $K$.
Note that the bound (10.1) with $|\alpha|=1$ corresponds to the mesh elements $K\in\mathcal{T}_{h}$ being shape regular (as in, e.g., [53, Equation 3.3]).
@@ 10.3 High-order interpolation
 $\text{ if }\,\,{\rm curl}\,v=0\,\,\text{ then }\,\,{\rm curl}\,(\mathcal{J}_{h}v)=0.$ (10.4)
@@
$\|v-\mathcal{J}_{h}v\|_{L^{2}(\Omega)}\leq C\bigg(\frac{h}{L}\bigg)^{\ell}\sum_{i=1}^{n}\sum_{j=0}^{\ell}L^{j}\Big(|v|_{H^{j}(\Omega_{i})}+(kh_{K})|k^{-1}{\rm curl}\,v|_{H^{j}(\Omega_{i})}\Big)$
@@
$\|k^{-1}{\rm curl}\,(v-\mathcal{J}_{h}v)\|_{L^{2}(\Omega)}\leq C\bigg(\frac{h}{L}\bigg)^{\ell}\sum_{i=1}^{n}\sum_{j=1}^{\ell}L^{j}|{\rm curl}\,v|_{H^{j}(\Omega_{i})}.$
@@
Points 1, 2, and 3 are proved in §11.1, §11.2, and §11.3, respectively. Regarding Point 3: we show in Lemma 11.5 below that $\max\{\gamma_{\rm dv}(P),\gamma_{\rm dv}(P^{*})\}\leq Ckh(1+kh)$, i.e., for the first inequality in (2.24) to hold, $kh$ must be sufficiently small. This condition is indeed weaker when $kL\gg 1$ than the condition $(kh)^{2p}C_{\rm sol}$ is sufficiently small” arising from the second inequality in (2.24), since $C_{\rm sol}\geq CkL$ (as recalled in Remark 1.6).
@@
Since $H_{0}({\rm curl}\,;\Omega)\supset C^{\infty}_{0}(\Omega)$ (by, e.g., [57, Equation 3.42]) and $C^{\infty}_{0}(\Omega)$ is dense in $L^{2}(\Omega)$, ${\cal H}=H_{0}({\rm curl}\,;\Omega)$ is dense in ${\cal V}=L^{2}(\Omega)$.
@@
With ${\cal H}=H_{0}({\rm curl}\,,\Omega)$, by, e.g., [57, Theorem 3.31],
@@
Let $w:=k^{-1}\zeta_{1}{\rm curl}\,v$ so that $k^{-1}{\rm curl}\,w=f$. Observe that $\mathop{{\rm div}}(\zeta_{1}^{-1}w)=0$ and $\mathop{{\rm div}}f=0$. With ${\rm div}_{T}$ the surface divergence on $\partial\Omega$, by, e.g., [57, Equation 3.52], on $\partial\Omega$,
@@
To prove that Part (ii) of Assumption 2.3 holds, we seek to apply Lemma 11.2 with $v=\Pi_{1}u$, $\zeta_{1}=\mu^{-1},$ and $\zeta_{2}=\mathcal{E}=\epsilon$. Since $C^{\infty}(\overline{D})$ is dense in $H^{s}(D)$ for all $s\in\mathbb{R}$ and $D\subset\mathbb{R}^{3}$ open (see, e.g., [50, Page 77]), $L^{2}(\Omega)$ is dense in $(H^{j}_{\rm pw}(\Omega))^{*}$ for $j\geq 1$, so the assumption that $\cal V$$^{*}$ is dense in $({\cal Z}^{j})^{*}=(Z^{j})^{*}$ for $j\geq 1$ is satisfied. The bound (2.10) then follows from (11.3) if $\Pi_{1}u$ is in $H_{0}({\rm curl}\,,\Omega)$ and satisfies $\mathop{{\rm div}}(\epsilon\Pi_{1}u)=0$. Recall from (2.8) that $\Pi_{0}:{\cal H}\to{\cal H}$, and thus also $\Pi_{1}:{\cal H}\to{\cal H}$. Thus, $u\in H_{0}({\rm curl}\,,\Omega)$ implies that $\Pi_{1}u\in H_{0}({\rm curl}\,,\Omega)$. For the zero-divergence condition, observe that, since gradients are always inside the kernel of $\rm curl$}\,,
@@
Thus, $\mathop{{\rm div}}\Pi_{1}^{*}v=0$ for all $v\in L^{2}(\Omega)$ (by, e.g., [57, Equation 3.33]). By (3.2), $\mathop{{\rm div}}(\epsilon\Pi_{1}u)=\mathop{{\rm div}}(\mathcal{E}\Pi_{1}u)=\mathop{{\rm div}}(\Pi_{1}^{*}\mathcal{E}\Pi_{1}u)$ so that $\mathop{{\rm div}}(\epsilon\Pi_{1}u)=0$ as required.
@@
By, e.g., [57, Theorem 3.41], the kernel of the curl operator in $H_{0}({\rm curl}\,,\Omega)$ equals $(\nabla H^{1}_{0}(\Omega))\oplus K_{N}(\Omega)$, where
@@
(the normal cohomology space); note that the dimension of $K_{N}(\Omega)$ equals the number of connected components of $\partial\Omega$ minus one; see, e.g., [57, Theorem 3.42]. Let $\Pi^{{\cal V}}_{\nabla H^{1}_{0}(\Omega)}$ and $\Pi^{{\cal V}}_{K_{N}(\Omega)}$ be the $\cal V$}-orthogonal projections onto $\nabla H^{1}_{0}(\Omega)$ and $K_{N}(\Omega)$, respectively, so that
@@
[…]
@@
We now apply Theorem 9.1 with $u=\nabla\phi$, $\zeta=I$, and $\kappa=m$ (note that this is allowed because $\Omega$ is $C^{m+1}$ with respect to the partition). Since $\phi\in H^{1}_{0}(\Omega)$ implies that $\nabla\phi\in H_{0}({\rm curl}\,;\Omega)$ (see, e.g., [57, Equation 3.60/§B.3]), the boundary condition on $u$ in Theorem 9.1 is satisfied and, for $\ell=0,\ldots,m-1$,
@@
$\hskip-28.45274pt\leq C\Big(\left\|\nabla\phi\right\|_{L^{2}(\Omega)}+\left\|\nabla w\right\|_{H^{j-1}_{{\rm pw},k}(\Omega)}\Big)$
$\hskip-28.45274pt\leq C\Big(\left\|\nabla\phi\right\|_{L^{2}(\Omega)}+\left\|\nabla w\right\|_{Z^{j}_{k}}\Big)=C\Big(\left\|\nabla\phi\right\|_{{\cal V}}+\left\|\nabla w\right\|_{{\cal Z}^{j}}\Big);$